# Continual Learning with Weight Interpolation

### 05 Apr 2024 v1

@@ 1 Introduction
Research goal. We propose a novel approach to continual learning that combines weight interpolation for better consolidation of the network capabilities before and after new tasks become available, with experience replay for enhanced robustness to catastrophic forgetting. Previous experiments on loss barriers in previous works were made mostly with the assumption that networks trained from two independent initializations are in two different local minima and have similar loss values [1, 20]. In the case of continual learning, this assumption cannot be met, as models are subject to forgetting [14] of previously seen data. In [30] weight averaging was proposed to mitigate catastrophic forgetting for pretrained models, however, parameter symmetries were not considered. Similarly, authors of [25] utilize weight interpolation to mitigate forgetting in BERT models, but they do not apply weight permutation. Pena et al. [38] propose a new weight interpolation method based on Sinkhorn differentiation, but continual learning is not their primary focus, and the scope of continuous learning experiments is very limited. Authors of [47] introduced a new interpolation method that could be used for models trained with disjoint data distributions, however, they do not carry out continual learning evaluation. In this work, we abuse the conjecture about low loss volume being convex modulo permutation symmetries [12], as each task will have separate data distribution and, consequently, different loss landscapes. However, in our theoretical analysis, we show what conditions should be met to increase the chances of finding good weight permutation and successful interpolation.
@@
Summary. In this work, we study the potential applications of recent findings from the field of weight interpolation in continual learning [10]. Based on recent weight interpolation techniques, we propose a remarkably simple continual learning algorithm that performs weight interpolation after each task to mitigate forgetting. We base our approach on widely used experience replay methods. Before training with new data from a new task we store network weights. The training with new data is carried out without any changes from standard replay algorithms. After training we utilize weights trained on the current task and stored old weights to perform permutation and then interpolation. The permutation step aligns units of both networks, while interpolations allow for better knowledge consolidation, compared to replay-based algorithms alone. We show that our method can reduce forgetting in several rehearsal-based methods.
@@ 6 Conclusion
Future works. Future work will focus on using CLeWI as a part of concept drift adaptation mechanisms [23] and on extending it for other computer vision tasks, such as object detection or continual segmentation.

### 09 Apr 2024 v2

@@ 1 Introduction
Previous experiments on loss barriers were made mostly with the assumption that networks trained from two independent initializations are in two different local minima and have similar loss values [1, 20]. In the case of continual learning, this assumption cannot be met, as models are subject to forgetting [14] of previously seen data. In [31] weight averaging was proposed to mitigate catastrophic forgetting for pretrained models, however, parameter symmetries were not considered. Similarly, authors of [25] utilize weight interpolation to mitigate forgetting in BERT models, but they do not apply weight permutation. Pena et al. [39] propose a new weight interpolation method based on Sinkhorn differentiation, but continual learning is not their primary focus, and the scope of continuous learning experiments is very limited. Authors of [48] introduced a new interpolation method that could be used for models trained with disjoint data distributions, however, they do not carry out continual learning evaluation.
Research goal. We propose a novel approach to continual learning that combines weight interpolation for better consolidation of the network capabilities before and after new tasks become available, with experience replay for enhanced robustness to catastrophic forgetting.
@@
Summary. In this work, we study the potential applications of recent findings from the field of weight interpolation in continual learning [10]. Based on recent weight interpolation techniques, we propose a remarkably simple continual learning algorithm that performs weight interpolation after each task to mitigate forgetting. In this work, we abuse the conjecture about low loss volume being convex modulo permutation symmetries [12], as each task will have separate data distribution and, consequently, different loss landscapes. However, in our theoretical analysis, we show what conditions should be met to increase the chances of finding good weight permutation and successful interpolation.
We base our approach on widely used experience replay methods. Before training with new data from a new task we store network weights. The training with new data is carried out without any changes from standard replay algorithms. After training we utilize weights trained on the current task and stored old weights to perform permutation and then interpolation. The permutation step aligns units of both networks, while interpolations allow for better knowledge consolidation, compared to replay-based algorithms alone. We show that our method can reduce forgetting in several rehearsal-based methods.
@@ 6 Conclusion
Future works. Future work will focus on exploring of the weight interpolation should be performed after every new task, or would a selective mechanism deciding when to perform interpolation lead to more robust results. Furthermore, we will explore the potential of using CLeWI as a part of concept drift adaptation mechanisms [23] and study the possibilities of extending it for other computer vision tasks, such as object detection or continual segmentation.
Acknowledgment
This work is supported by the CEUS-UNISONO programme, which has received funding from the National Science Centre, Poland under grant agreement No. 2020/02/Y/ST6/00037.
@@ 10 Memory restricted evaluation of rehearsal methods
11 Implementation
The CLeWI algorithm is straightforward to implement and can be used as a plug-in for existing continual learning algorithms. This is in line with the construction of modern continual learning frameworks such as Mammoth [5], or Avalanche [29]. In such a case, one needs to implement only a method that is called after training with a given task that performs interpolation and stores the previous model. This should allow for easy combining of existing methods for improved performance. However, one needs to be careful during implementation. Other algorithms could also take some actions after training with a given task. Whether these actions are taken before or after the interpolation could impact the overall algorithm performance. This should be considered case-by-case, as different algorithms could require different orders of post-task training actions. In many cases, this can be trivial, as other algorithms do not collide with weight interpolation, but there are a few that require extra steps in the implementation. For example, we found that in the case of BiC [53], first, we need to compute the bias correction, then store the old model, then update the buffer, perform weight interpolation, and evaluate bias compute bias correction for interpolated models again. As a result, for BiC+CLeWI, we keep three copies of models instead of two: one for the current model, the second for BiC knowledge distillation, and the last for weight interpolation. Using the same model for weight interpolation and knowledge distillation provided worse results. The model before interpolation has better representation of most recent task, therefore could be more suitable for knowledge destination. This should be taken into consideration when combining CLeWI with other algorithms that store weights our utilize activations of a network trained on previous tasks (such as LwF [28]). Combining CLeWI with other methods that require post-training actions may require similar modifications.